NPP Projection Close 11/25
To do this evening/ was completed:
Presentation Issues: Scope masking 2.39 cue was built into the Spartan SPL but was not wide enough for the picture...Lillian was in the auditorium watching and said to trigger the film masking cue but I did not see her message in time...she came up to 4/6 and hit the cue on the quick controls...the image looked fine for the rest of the screening
Preshow Clips/Playlist: Converted the intro slide for Nightmare on Elm Street 4 to DCP and saved it to the P&T drive...later ingested it to the LMS
Trailer Updates:
Special Events/ AV set up: Set up the mic and mixer in theater 4 for the intro to Spartan...intro and screening went smoothly (changeovers were nice and clean) save for a brief interlude when the lamp in projector 2 began to go dark and back to light...realized I'd had the damn exhaust valves turned the wrong way thinking they were open (got too used to the valves on booth 1's machines)...and brought the house lights up and corrected this with Lillian's assistance
Presentation Issues: Scope masking 2.39 cue was built into the Spartan SPL but was not wide enough for the picture...Lillian was in the auditorium watching and said to trigger the film masking cue but I did not see her message in time...she came up to 4/6 and hit the cue on the quick controls...the image looked fine for the rest of the screening
Preshow Clips/Playlist: Converted the intro slide for Nightmare on Elm Street 4 to DCP and saved it to the P&T drive...later ingested it to the LMS
Trailer Updates:
Special Events/ AV set up: Set up the mic and mixer in theater 4 for the intro to Spartan...intro and screening went smoothly (changeovers were nice and clean) save for a brief interlude when the lamp in projector 2 began to go dark and back to light...realized I'd had the damn exhaust valves turned the wrong way thinking they were open (got too used to the valves on booth 1's machines)...and brought the house lights up and corrected this with Lillian's assistance
Cleaned both projectors and returned the mic/mixer equipment back to its proper storage
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Transferred Holy Grail and Life of Brian from George Ivan to the LMS
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Transferred Holy Grail and Life of Brian from George Ivan to the LMS
2000 Meters finished transferring from Aspera
Outgoing/Shipping (content):
Print/video Inspection: Spartan is tails out in booth 4
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Double checked everything for tomorrow...had to stack SPLs for Hamnet and Eternity in the penultimate and final rounds in 5 since the flip time was so short so keep an ear out for any lights up requests and/or if playback fails and you have to skip ahead in the latter show
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?: Yes
Projection cart by entrance?: Yes
Additional Notes:
Outgoing/Shipping (content):
Print/video Inspection: Spartan is tails out in booth 4
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Double checked everything for tomorrow...had to stack SPLs for Hamnet and Eternity in the penultimate and final rounds in 5 since the flip time was so short so keep an ear out for any lights up requests and/or if playback fails and you have to skip ahead in the latter show
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?: Yes
Projection cart by entrance?: Yes
Additional Notes:
Comments
Post a Comment